OpenStudy (anonymous):

find the quotient -42s^3+6s^2+64s/2s

OpenStudy (campbell_st):

same process \[\frac{2 \times -21 s \times s \times s}{2 \times s} + \frac{2 \times 3 \times s \times s}{2 \times 2} + \frac{2 \times 32 \times s}{2 \times s}\] what are you left with..?

OpenStudy (anonymous):

-21s and s 2, 3, s, s 32

OpenStudy (campbell_st):

so -21s^2 +3s + 32